OpenStudy (anonymous):

use the square root property to solve the quadratic equation (x-3)^2=4 a.x=7 b.x=1,-5 c. x=5,1 d. x=2,-2

OpenStudy (anonymous):

First of all: (x-3)^2=4 \[\sqrt{(x-3)^{2}}= \pm \sqrt{4}\] \[x-3 = \pm 2 \] \[\left\{ 1,5 \right\}\]

OpenStudy (anonymous):

Taking the square root of \[(x - 3)^{2}\] saves you a lot of work since 4 is a perfect square. This method is great for solving exercises like this one.

OpenStudy (anonymous):

thanks :)